# Zeroth-Order Policy Gradient for Reinforcement Learning from Human Feedback without Reward Inference

### v1

@@ 5.3 Technical Challenges and Proof Novelties
Classic Proof of Zeroth-Order Optimization: to illustrate the procedure of the analysis of zeroth order gradient estimate, we suppose we can query $V(\pi_{{\bm{\theta}}})$ for any ${\bm{\theta}}$. This procedure makes use of the randomized smoothing function $V_{\mu}({\bm{\theta}})$ (Ghadimi and Lan, 2013; Gao et al., 2018):
@@
The standard proof uses the randomized smoothing function $V_{\mu}(\pi_{{\bm{\theta}}})$ as the Lyapunov function and then bounds the drift given the stochastic gradient descent update rule when $\alpha=\Theta(1/d)$. Neglecting problem independent constants, we have:
@@
$\leq$ $-\alpha\underbrace{\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}}_{\mathsf{NegDrift}}+\alpha\underbrace{\langle\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}}),\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})-\hat{{\bm{g}}}_{t}}_{\text{$1$st Order: }\mathsf{GradBias}}\rangle+\alpha^{2}\underbrace{\|\hat{{\bm{g}}}_{t}-\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}}_{\text{2nd Order: }\mathsf{GradVar}\approx\mu^{2}d^{2}}.$
@@
$\underbrace{\frac{\mathbb{E}[V_{\mu}(\pi_{{\bm{\theta}}_{0}})]-\mathbb{E}[V_{\mu}(\pi_{{\bm{\theta}}^{*}})]}{T}}_{{\mathcal{O}}(H/T)}\lesssim-\alpha\underbrace{\mathbb{E}\left[\frac{\sum_{t=1}^{T}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}}{T}\right]}_{\mathsf{Target}}+\alpha\mu^{2}d^{2}.$
@@
Amplified Gradient Biases for ZPG: If we directly apply the steps above to ZPG, we immediately run into the issue that our gradient estimator $\hat{{\bm{g}}}_{t}$ in expectation is biased even compared to smoothing function gradient due to preference estimation. Moreover, the gradient variance on the second order term will also be larger since we used trajectory reward to estimate the value function. Through concentration inequalities, we will be able to obtain a bound on the error of using preference to estimate the value function difference as:
$\left|\mathbb{E}\left[\sigma^{-1}(p_{t,n})\right]-\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})\right)\right|\leq\tilde{{\mathcal{O}}}\left(1/\sqrt{M}\right),$
@@
Implicit Smoothing Function for ZBCPG: Due to the choice of blocked perturbation vector sampling procedure, it is difficult to obtain the exact analytical expression of the smoothing function $V_{\mu}(\pi_{{\bm{\theta}}})$ whose gradient is the expectation of gradient estimation $\hat{{\bm{g}}}_{t}$ for ZBCPG. This prohibits us from continuing to use $V_{\mu}(\pi_{{\bm{\theta}}})$ as the Lyapunov function, as it is hard to analyze the gradient bias and the variance without an explicit target format. However, if we rethink why we introduce the smoothed function in zeroth-order optimization, what we hope for is the gradient of the smoothed function will be unbiased to cancel out the first order positive drift. However, this is already not true in the analysis of ZPG since we have gradient estimation bias from human feedback, but it is small enough on average and we can control it. If the gradient difference between the smoothed function $V_{\mu}(\pi_{{\bm{\theta}}})$ and the vanilla value function $V(\pi_{{\bm{\theta}}})$ is smaller than this additional bias, then we can use the original value function $V(\pi_{{\bm{\theta}}})$ as the Lyapunov function at the cost of an additional bias besides preference estimation. Fortunately, this can be achieved through a carefully chosen perturbation distance $\mu$ to balance these two errors.
@@ 6 Conclusion
In this paper, we proposed two RLHF algorithms without reward inference based on zeroth-order policy gradient called ZPG and ZBCPG, which train the policy network directly from human preferences without a global reward function approximator. Both algorithms are shown to have a provable polynomial sample complexity to learn a stationary policy under mild conditions, which shows there exist provably efficient methods to solve RLHF problems in general MDPs without reward inference.
Appendix A Proof Backbone for Theorems: Preference Estimation Error
@@
This lemma shows that our reward difference estimation from human preference is accurate as long as the number of human evaluators $M$ is large. The proof of this lemma is provided in Sec. A.1
A.1 Proof of Lemma. 1
@@
$\leq$ $\exp\left(-2M\frac{\log\frac{1}{\delta}}{M}\right)$
$\leq$ $\delta,$
@@
$\left|\sigma^{-1}(p_{t,n})-[r(\tau_{n,1})-r(\tau_{n,0})]\right|=$ $\left|\sigma^{-1}(p_{t,n})-\sigma^{-1}({\mathbb{P}}(\tau_{n,1}\succ\tau_{n,0}))]\right|$
$\leq$ $L\left|p_{t,n})-{\mathbb{P}}(\tau_{n,1}\succ\tau_{n,0}))\right|$
@@
where the last inequality uses Lemma. 2. Combining both cases and let $\delta=1/M^{2}$, we have:
@@
When the concentration holds, we also employ the continuity of the inverse link function as:
@@
Appendix B Proof of Theorem. 1
@@
    $\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}})=\mathbb{E}_{{\bm{v}}}\left[\frac{d}{\mu}\left(V(\pi_{{\bm{\theta}}+\mu{\bm{v}}})-V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}})\right){\bm{v}}\right].$ (7)
@@
$\mathbb{E}\left[V_{\mu}(\pi_{{\bm{\theta}}_{t}})-V_{\mu}(\pi_{{\bm{\theta}}_{t+1}})|{\mathcal{F}}_{t}\right]\leq$ $-\frac{\alpha}{2}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}+\frac{\alpha\mu^{2}L^{2}d^{2}}{4}+\frac{L\alpha^{2}}{2}\underbrace{\mathbb{E}\left[\left.\|\hat{{\bm{g}}}_{t}\|_{2}^{2}\right|{\mathcal{F}}_{t}\right]}_{\mathsf{Var}_{t}}$ (11)
$+\alpha\underbrace{\langle\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}}),\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})-\mathbb{E}\left[\left.\hat{{\bm{g}}}_{t}\right|{\mathcal{F}}_{t}\right]\rangle}_{\mathsf{Bias}_{t}}.$
Then, we analyze the two positive drift terms $\mathsf{Bias}_{t}$ for the first-order gradient bias and $\mathsf{Var}_{t}$ for the gradient noise separately, and then construct a tight upper bound of the Lyapunov drift. The results are summarized in the following two lemmas where the proofs are deferred to Sec. B.1 and Sec. B.2:
@@
$\leq$ $\frac{2dL}{\mu}\sqrt{\frac{\log M}{M}}\left\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\right\|_{2}+d^{2}L^{2}\sqrt{\frac{\log M}{M}}.$
@@
we have the positive first-order drift regarding gradient norm is bounded by:
$\frac{2\alpha dL}{\mu}\sqrt{\frac{\log M}{M}}\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2}\leq\frac{1}{8}\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2},$
@@
If on the other hand, the gradient is large, then we can upper bound the first-order drift regarding gradient norm in another way:
@@
$\leq$ $-\frac{\alpha}{4}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}+\frac{\alpha\mu^{2}L^{2}d^{2}}{2}+\frac{32\alpha d^{2}L^{2}\log M}{\mu^{2}M}+\alpha d^{2}L^{2}\sqrt{\frac{\log M}{M}}$

### v2

@@ 4.3 Technical Challenges and Proof Novelties
Classic Proof of Zeroth-Order Optimization: to illustrate the procedure of the analysis of zeroth order gradient estimate, we suppose we can query $V(\pi_{{\bm{\theta}}})$ for any ${\bm{\theta}}$. This procedure makes use of the randomized smoothing function $V_{\mu}({\bm{\theta}})$ (Ghadimi and Lan, 2013; Gao et al., 2018) as
@@
The standard proof uses the randomized smoothing function $V_{\mu}(\pi_{{\bm{\theta}}})$ as the Lyapunov function and then bounds the drift given the stochastic gradient descent update rule when $\alpha=\Theta(1/d)$. Neglecting problem-independent constants, we have:
@@
$\leq$ $-\alpha\underbrace{\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}}_{\mathsf{Drift}}+\alpha\underbrace{\langle\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}}),\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})-\hat{{\bm{g}}}_{t}}_{\text{$1$st Order: }\mathsf{GradBias}}\rangle+\alpha^{2}\underbrace{\|\hat{{\bm{g}}}_{t}-\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}}_{\text{2nd Order: }\mathsf{GradVar}\approx\mu^{2}d^{2}}.$
@@
$\underbrace{\frac{\mathbb{E}[V_{\mu}(\pi_{{\bm{\theta}}_{0}})]-\mathbb{E}[V_{\mu}(\pi_{{\bm{\theta}}_{T}})]}{T}}_{{\mathcal{O}}(H/T)}\lesssim-\alpha\underbrace{\mathbb{E}\left[\frac{\sum_{t=1}^{T}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}}{T}\right]}_{\mathsf{Target}}+\alpha\mu^{2}d^{2}.$
@@
Amplified Gradient Biases for ZPG: If we directly apply the steps above to ZPG, we immediately run into the issue that our gradient estimator $\hat{{\bm{g}}}_{t}$ in expectation is biased even compared to smoothing function gradient due to preference estimation. Moreover, the second-order gradient variance will be larger since we used trajectory reward to estimate the value function. From concentration, we will obtain an error bound of using preference to estimate the value function difference as:
$\left|\mathbb{E}\left[\sigma^{-1}(p_{t,n})\right]-\left(V(\pi_{{\bm{\theta}}_{t}+\mu{\bm{v}}_{t}})-V(\pi_{{\bm{\theta}}_{t}})\right)\right|\leq\tilde{{\mathcal{O}}}\left(\frac{1}{\sqrt{M}}\right),$
@@
Implicit Smoothing Function for ZBCPG: Due to the choice of blocked perturbation vector sampling procedure, it is difficult to obtain the exact analytical expression of the smoothing function $V_{\mu}(\pi_{{\bm{\theta}}})$ whose gradient is the expectation of gradient estimation $\hat{{\bm{g}}}_{t}$ for ZBCPG. This prohibits us from continuing to use $V_{\mu}(\pi_{{\bm{\theta}}})$ as the Lyapunov function, as it is hard to analyze the gradient bias and the variance without an explicit target format. However, if we rethink the reason for introducing the smoothed function in zeroth-order optimization, we hope the gradient of the smoothed function will be unbiased to cancel out the first-order positive drift. However, this is already not true in the analysis of ZPG since we have gradient estimation bias from human feedback, but it is small enough on average to be controlled. If the gradient difference between the smoothed function $V_{\mu}(\pi_{{\bm{\theta}}})$ and the vanilla value function $V(\pi_{{\bm{\theta}}})$ is smaller than this additional bias, then we can use the original value function $V(\pi_{{\bm{\theta}}})$ as the Lyapunov function at the cost of an additional bias besides preference estimation. Fortunately, this can be achieved through a carefully chosen perturbation distance $\mu$ to balance these two types of errors.
5 Experiments
In this section, we study the empirical performance of ZPG and ZBCPG in a stochastic GridWorld environment to show the practical value of both algorithms, which coincides with the theoretical results. We consider three baselines: (1) RM+PPO (Ouyang et al., 2022), (2) DPO for token-level MDP (Rafailov et al., 2024b), and (3) Online DPO for token-level MDP (Dong et al., 2024; Guo et al., 2024), where the implementation detail can be found in the appendix.
Environment. Our GridWorld environment consists of $5\times 5$ blocks, where the location of each block, i.e., the state space ${\mathbb{S}}$, is denoted as $(1,1)$ to $(5,5)$. For each block, we flip a fair coin. If the head shows, we will place a random reward on top, and if the tail shows, there will be no reward on this block. The reward, if placed in a block, is randomly sampled from a standard normal distribution, which means we allow a negative reward as punishment. Each episode consists of $H=20$ steps, and at the start of each episode, an agent is positioned in block $(3,3)$, i.e., the center of the GridWorld environment. At each step, the agent can choose from four actions that constitute the action space ${\mathbb{A}}$: going up, going down, going left, and going right. However, unlike classic GridWorld environments where the agent can perfectly control its movement, in our experiment, we assume the agent is subject to environment disturbances or imperfect control such as wind or turbulence, to model the stochastic transition and demonstrate the weaknesses of predominant RLHF frameworks such as DPO. We assume at each step, the action taken by the agent may be reversed with probability $0.4$ due to disturbances. For example, if the agent is in the center and chooses to go up, it will land on the block above the current block with a probability of $0.6$ and on the block beneath it with a probability of $0.4$. Similarly, if the agent chooses to go left, it will land on the block left next to it with a probability of $0.6$ and on the block right next to it with a probability of $0.4$. The motivation for imperfect control arises naturally from designing agents for turbulent environments, e.g., for robots operating on airplanes, cars, or ships, where continual wind or bump may shift the agent’s action. The goal of the agent is to maximize the cumulative reward over the $H=20$ time horizon, and the interaction is conducted episodically.
Policy Parameterization. For all algorithms implemented in our experiment, we used tabular policy softmax parameterization, i.e., each state-action pair $(s,a)$ is equipped with a parameter $\xi_{s,a}$ and the policy $\pi(a|s)$ of taking action $a$ at state $s$ would follow:
$\pi(a|s)=\frac{\exp(\xi_{s,a})}{\sum_{a}\exp(\xi_{s,a})}.$
Human Feedback. In this experiment, we assume access to a human panel of size $M=1000$, i.e., a group of human experts, can be recruited and queried at the same time for two trajectories, except for the distributed implementation of ZBCPG, which we will discuss in more detail later. Each human expert, which we refer to as a panelist, will generate a preference based on a specific preference model characterized by a link function. The agent can query human evaluators for preference over two trajectories. We consider two preference models: (1) Bradley-Terry model with a logistic link function and (2) Weibull model with an anti-asymmetric Weibull link function, both over the trajectory rewards. For example in Bradley-Terry model, for two trajectories $\tau_{0}$ and $\tau_{1}$, each panelist would provide feedback following the distribution as follows:
$\mathbb{P}(\tau_{1}\succ\tau_{0})=\frac{\exp(r(\tau_{1}))}{\exp(r(\tau_{1}))+\exp(r(\tau_{0}))},\quad\mathbb{P}(\tau_{0}\succ\tau_{1})=\frac{\exp(r(\tau_{0}))}{\exp(r(\tau_{1}))+\exp(r(\tau_{0}))}.$
For the Weibull preference feedback mechanism (Train, 2009), it is worth noting that the link function is anti-asymmetric between the two trajectories being compared, i.e., $\sigma(x)\neq 1-\sigma(-x)$. For example, if two trajectories $(\tau_{0},\tau_{1})$ are compared by a panelist with the Weibull preference model, it would prefer the trajectory $\tau_{1}$ with probability:
$\mathbb{P}(\tau_{1}\succ\tau_{0})=\exp_{2}(-\exp_{2}(r(\tau_{0})-r(\tau_{1}))).$
This preference model would capture the bias effect presented in the revealing sequence to the human evaluators. Namely, human experts would unintentionally find it easier to distinguish if the better trajectory is presented first.
5.1 Results on Bradley-Terry Human Feedback
(a) performance
(b) parallelization
Figure 2: GridWorld with Bradley-Terry Feedback: (a) the trajectory return of ZPG, ZBCPG, and RLHF baselines, and (b) the return of ZBCPG with different parallelization levels. All results are averaged over $10^{5}$ repetitions of policy evaluation and shaded areas indicate confidence intervals.
Algorithm ZPG ( Ours) ZBCPG ( Ours) RM+PPO DPO Online DPO
Return $\bm{1.94}\pm 0.09$ $\bm{1.91}\pm 0.09$ $1.80\pm 0.09$ $-4.13\pm 0.09$ $1.71\pm 0.09$
Table 1: Last Iterate Policy Average Return with Bradley-Terry Feedback.
Comparison of Average Return. We first study and compare the empirical performance of all algorithms with Bradley-Terry human feedback. All algorithms, including both our proposed ZPG and ZBCPG with all baselines, collect $N=1000$ trajectory pairs between policy updates and $M=1000$ human experts evaluate each pair. The trajectory return for each iteration is compared in Fig. 2(a) and the return of the final policy is reported in Tab. 1. Both ZPG and ZBCPG perform better than the three baselines in both convergence rates and the quality of the last iterate policy. Compared to PPO, our algorithms converge to a better policy partially since the reward model is inaccurate and the agent is not able to learn the optimal policy from it. It is also observed that vanilla DPO has a much worse performance than our proposed algorithms. This may result from two reasons: first, the DPO loss is valid only in deterministic MDP, and second, DPO is constrained to the neighborhood of the sub-optimal reference policy. The online DPO algorithm improves over vanilla DPO but still has inferior performance due to the inherent model error of the DPO loss. This also shows the fundamental difference between stochastic and deterministic MDPs, and the need to design RLHF algorithms for general RL problems. Moreover, online DPO converges much slower, partly because the DPO loss landscape becomes flat and hard to optimize when the weight of the KL constraint is small for better exploration.
Power of Parallelization. We also compare our proposed algorithms to distributed implementations of ZBCPG to show the influence of parallelization of the block-coordinate descent. For distributed ZBCPG, we assume multiple parallel agents can collect trajectories at the same time and query humans at the same time. However, the panel of human evaluators is also separated into small groups for parallelization and each agent can only work with a small group of experts to perform optimization. The results are shown in Fig. 2(b) where the number of blocks shown in the legend controls the level of parallelization. A detailed description of the implementation is presented in the appendix. It is shown that as the number of blocks increases, ZBCPG converges faster to a stationary policy. However, the number of human queries per pair of trajectories in each parallelization also decreases, which introduces a larger gradient bias and leads to a sub-optimal policy. Therefore, the trade-off between computation parallelization and accuracy should be taken carefully.
5.2 Results on Weibull Human Feedback
In this section, we conduct additional numerical experiments on Weibull human feedback to evaluate the impact of the preference model link functions on our proposed algorithms and the baselines, as well as the influence of the size $M$ of the human evaluator panel.
(a) Weibull, $M=1000$
(b) Weibull, $M=200$
Figure 3: GridWorld with Weibull Feedback: (a) the return of ZPG, ZBCPG, and RLHF baselines under Weibull human feedback with panel size $M=1000$, and (b) the trajectory return of ZPG, ZBCPG, and RLHF baselines under Weibull human feedback with panel size $M=200$. All results are averaged over $10^{5}$ repetitions of policy evaluation and shaded areas indicate confidence intervals.
Influence to Return Comparison. The implementations of all algorithms are the same as in Sec. 5 except that we use the Weibull link function in both ZPG and ZBCPG. The results are shown in Fig. 3(a) and Tab. 2. It can be seen that both our algorithms perform much better than the baselines in terms of convergence rate and the quality of the last iterate policy. Even though the trend is similar to the Bradley-Terry feedback results shown in Sec. 5, the performance gap is slightly larger due to the additional model mismatch factor. Both DPO and Online DPO are designed explicitly under the Bradley-Terry model, which is inaccurate for the ground-truth Weibull feedback, and the same applies to the reward model training step in the PPO baseline. However, the performance is not significantly degrading, which implies the Weibull link function is not fundamentally different than the Bradley-Terry model and allows these algorithms to have a compelling performance still. It is also shown that our algorithms are more flexible with the different human preference models and deliver better quality.
Algorithm ZPG ( Ours) ZBCPG ( Ours) RM+PPO DPO Online DPO
Return $\bm{2.13}\pm 0.09$ $\bm{1.90}\pm 0.09$ $1.72\pm 0.09$ $-4.25\pm 0.09$ $1.69\pm 0.09$
Table 2: Last Iterate Policy Average Return under Weibull Feedback with $M=1000$.
Panel Size. To study the influence of panel size $M$ on the quality of the learned policy, and to compare with our theoretical results presented in Theorem. 1 and Theorem 2, we conduct the same experiment with Weibull preference feedback with the same setup as the previous subsection where we only use $M=100$ human evaluators for each trajectory pair, which is smaller than the ${\mathcal{O}}(H^{2})$ guarantee required by theory. The results are shown in Fig. 3(b) and Tab. 3. It can be seen that even with a much smaller human panel, our proposed algorithms can still converge to a local optimal policy which has a better quality than all benchmarks. Comparing the quality of the last iterate policy in Tab. 3 to Tab. 2, it can also be deduced that the final policy found by our proposed algorithm is robust to the size of the human panel $M$, and the performance with less human evaluators does not significantly degenerate. However, it can also be seen from Fig. 3 that both ZPG and ZBCPG have relatively larger policy quality fluctuation over the training trajectory compared to the three baselines, and this fluctuation increases as the number of human experts $M$ decreases, which poses stability concerns. This is partially because our proposed algorithms do not use KL regularization over the policies to constrain each gradient ascent step, which results in large policy alterations between iterations. We conjure that this is likely to be mitigated if we apply the KL regularization to our algorithms.
Algorithm ZPG ( Ours) ZBCPG ( Ours) RM+PPO DPO Online DPO
Return $\bm{2.02}\pm 0.08$ $\bm{2.16}\pm 0.09$ $1.63\pm 0.09$ $-4.22\pm 0.10$ $1.69\pm 0.09$
Table 3: Last Iterate Policy Average Return under Weibull Feedback with $M=100$.
@@ 6 Conclusion
In this paper, we proposed two RLHF algorithms without reward inference based on a zeroth-order policy gradient called ZPG and ZBCPG. Both algorithms are shown to have a provable polynomial sample complexity to learn a stationary policy under mild conditions and exhibit nice empirical performances in environments with stochastic transitions, outperforming popular RLHF baselines.
Limitations and Future Directions. Both proposed algorithms require many online human queries for accurate gradient estimation to ensure fast convergence to a good policy, both in theory and in experiments, which may limit their practicality. Such limitation is common in all online RLHF algorithms, including online DPO (Guo et al., 2024). On the other hand, online RLHF algorithms exhibit better performances, which is also observed in our experiments, which may converge faster and require fewer iterations. One possible approach to avoid this limitation is to replace human feedback with AI feedback similar to Guo et al. (2024). Another possible method is to replace the SGD style update in the current algorithm with a momentum-based optimization scheme, to reuse the human queries from previous policy iterations. The necessity of the current level of human queries will be further explored. Another limitation of our algorithms is the lack of strategic exploration, which may either fail to explore hard-to-reach trajectories or produce similar trajectories that are hard to provide preferences. One possible way to incorporate better exploration practice into our proposed algorithms is to replace the SGD type of policy update with projected SGD, which forces the agent to explore before convergence. Solving both limitations will be our future direction.
Acknowledgments
The work of Qining Zhang and Lei Ying is supported in part by NSF under grants 2112471, 2134081, 2207548, 2240981, and 2331780.
Appendix A Details of the Numerical Experiment in Sec. 5
In this section, we describe the implementation of our proposed algorithms and the implementation of baseline algorithms such as DPO, online DPO, and PPO.
A.1 Experiments Comparing ZPG and ZPBCPG to Baselines
In this section, we describe the implementation of algorithms in Fig. 2(a), Fig. 3(a) and (b). For all algorithms, we tune the learning rates and perturbation distances if necessary. We also ensure the total number of trajectory samples and the total number of human queries are the same for fair comparison.
ZPG. We implemented ZPG according to Alg. 1 with trim size $\Delta=0.001$ and batch size $N=1000$ for $T=1000$ iterations. After collecting $N$ trajectory pairs for the perturbed policy and the current policy, all panelists were queried to obtain the preference of all trajectory pairs.
ZBCPG. We implemented ZBCPG according to Alg. 2 with the same setup as ZPG with trim size $\Delta=0.001$ and batch size $N=1000$ for $T=1000$ iterations. The coordinate batch size $K$ is chosen to be $5\times 4=20$, i.e., at each time we perturb the parameter $\xi_{s,a}$ associated with all actions for $5$ states simultaneously. After collecting $N=1000$ trajectory pairs for the perturbed policy and the current policy, all panelists were queried to obtain the preference of all trajectory pairs.
RM+PPO. The PPO baseline contains two components: reward model training and policy training. The reward model we consider used tabular parameterization. To train the reward model, we first used the randomly initialized policy to collect $5\times 10^{5}$ trajectory pairs, which is half the number of trajectories used in ZPG and ZBCPG. Then, we queried all panelists for the preferences of all trajectory pairs and obtained the population preference probability for all pairs. After that, we used SGD optimizer to train the reward model for $5$ epochs under the loss in Christiano et al. (2017), which also used multiple human queries for each pair of trajectories. For the policy training, we used standard online PPO with a KL regularization weight being $0.1$. We also conducted $T=1000$ policy iterations, and the agent collects $N=1000$ trajectories between policy updates for improved training stability. The PPO loss in each iteration is optimized using SGD for $5$ epochs.
DPO. To train the DPO agent, we also conducted $T=1000$ policy iterations, and the agent collects $N=1000$ trajectory pairs between policy updates. The KL regularization weight in the DPO loss is $0.1$. For each pair of trajectories, the agent queried all panelists for preferences and then used the averaged preference probability to replace the single human preference in the vanilla DPO loss, similar to the loss in Christiano et al. (2017), to ensure fair comparison among algorithms. The DPO loss at each iteration is optimized using SGD for $5$ epochs.
Online DPO. The implementation of online DPO is almost the same as offline DPO, except that we replaced the reference policy in the KL regularization with the current policy.
A.2 Experiment Comparing ZPG to Distributed ZBCPG Implementation
In this section, we describe the distributed implementation of ZBCPG, which produced the results in Fig. 2(b). For distributed ZBCPG, we assume multiple parallel agents can collect trajectories and query humans in parallel, and the number of blocks in the legend of Fig. 2(b) controls the level of parallelization. For example, when the number of blocks is $5$, we divide the parameters in the policy network into $5$ blocks. Then, five perturbed policies are created by perturbing each block individually. These policies are distributed to $5$ agents, and each agent will collect $N=1000$ trajectories simultaneously. However, to ensure fairness in the human panel query, we also divide the complete $M=1000$ people panel into $5$ sub-panels that could work in parallel, where each sub-panel includes $200$ panelists and will assist one agent with human preference queries. Therefore, the parallelized version of ZBCPG only has the advantage of sampling multiple trajectories, but not the access to more human experts. When the number of blocks is $5$, we perturbed the parameters associated with all actions for $5$ states simultaneously so each block contains $5\times 4=20$ parameters. Then, the $5$ blocks are distributed to $5$ agents to estimate the policy gradient. When the number of blocks is $25$, we perturbed parameters associated with all actions for a single state and each block contains $4$ parameters. When the number of blocks is $1$, we perturbed all parameters with Rademacher noise. The implementation of ZPG and all other parameters are the same as in the previous section.
Appendix B Proof Backbone for Theorems: Preference Estimation Error
@@
This lemma shows that our reward difference estimation from human preference is accurate as long as the number of human evaluators $M$ is large. The proof of this lemma is provided in Sec. B.1
B.1 Proof of Lemma. 1
@@
$\leq$ $\exp\left(-2M\frac{\log\frac{1}{\delta}}{M}\right)\leq\delta,$
@@
$\left|\sigma^{-1}(p_{t,n})-[r(\tau_{n,1})-r(\tau_{n,0})]\right|=$ $\left|\sigma^{-1}(p_{t,n})-\sigma^{-1}({\mathbb{P}}(\tau_{n,1}\succ\tau_{n,0}))\right|$
$\leq$ $L\left|p_{t,n}-{\mathbb{P}}(\tau_{n,1}\succ\tau_{n,0}))\right|$
@@
where the last inequality uses the smoothness of the inverse link function, i.e., Lemma. 2. Combining both cases and let $\delta=1/M^{2}$, we have:
@@
This is because the inverse link function is bounded after applying the trim operator to the preference probability estimation. When the concentration holds, we also employ the continuity of the inverse link function as:
@@
Appendix C Proof of Theorem. 1
@@
    $\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}})=\mathbb{E}_{{\bm{v}}}\left[\frac{d}{\mu}\left(V(\pi_{{\bm{\theta}}+\mu{\bm{v}}})-V(\pi_{{\bm{\theta}}})\right){\bm{v}}\right].$ (7)
@@
$\mathbb{E}\left[V_{\mu}(\pi_{{\bm{\theta}}_{t}})-V_{\mu}(\pi_{{\bm{\theta}}_{t+1}})|{\mathcal{F}}_{t}\right]$ (11)
$\leq$ $-\frac{\alpha}{2}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}+\frac{\alpha\mu^{2}L^{2}d^{2}}{4}+\frac{L\alpha^{2}}{2}\underbrace{\mathbb{E}\left[\left.\|\hat{{\bm{g}}}_{t}\|_{2}^{2}\right|{\mathcal{F}}_{t}\right]}_{\mathsf{Var}_{t}}+\alpha\underbrace{\langle\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}}),\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})-\mathbb{E}\left[\left.\hat{{\bm{g}}}_{t}\right|{\mathcal{F}}_{t}\right]\rangle}_{\mathsf{Bias}_{t}}.$
Then, we analyze the two positive drift terms $\mathsf{Bias}_{t}$ for the first-order gradient bias and $\mathsf{Var}_{t}$ for the gradient noise separately and then construct a tight upper bound of the Lyapunov drift. The results are summarized in the following two lemmas where the proofs are deferred to Sec. C.1 and Sec. C.2:
@@
$\mathsf{Bias}_{t}=$ $\langle\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}}),\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})-\mathbb{E}\left[\left.\hat{{\bm{g}}}_{t}\right|{\mathcal{F}}_{t}\right]\rangle\leq\frac{2dL}{\mu}\sqrt{\frac{\log M}{M}}\left\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\right\|_{2}+d^{2}L^{2}\sqrt{\frac{\log M}{M}}.$
@@
we have the positive first-order drift regarding the gradient norm is bounded by:
$\frac{2\alpha dL}{\mu}\sqrt{\frac{\log M}{M}}\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2}\leq\frac{\alpha}{8}\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2}^{2},$
@@
If on the other hand, the gradient is small, i.e., we have:
$\left\|\nabla_{{\bm{\theta}}}V_{\mu}(\pi_{{\bm{\theta}}_{t}})\right\|_{2}\leq\frac{16dL}{\mu}\sqrt{\frac{\log M}{M}},$
and then we can upper bound the first-order drift regarding gradient norm in another way:
@@
$\leq$ $-\frac{\alpha}{4}\|\nabla_{{\bm{\theta}}}V(\pi_{{\bm{\theta}}_{t}})\|_{2}^{2}+\frac{\alpha\mu^{2}L^{2}d^{2}}{2}+\frac{32\alpha d^{2}L^{2}\log M}{\mu^{2}M}+\alpha d^{2}L^{2}\sqrt{\frac{\log M}{M}}+\frac{5\alpha^{2}d^{2}L^{3}\log M}{\mu^{2}M}+\frac{6\alpha^{2}d^{2}H^{2}L}{\mu^{2}N}$